OpenStudy (anonymous):

Which equation is equivalent to 2x + 6y = 12?

OpenStudy (anonymous):

out of other possibilities?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

A. y=1/3x+2

OpenStudy (anonymous):

b. y=1/3x-2

OpenStudy (anonymous):

c. y=-1/3x+2

OpenStudy (anonymous):

d. y=-1/3x-2

OpenStudy (anonymous.):

I will go with "A"

OpenStudy (anonymous):

if its right i will give u a medal (:

OpenStudy (anonymous.):

Wait .. :3

OpenStudy (anonymous.):

its "b"

OpenStudy (anonymous):

okay